# Bigger, Regularized, Optimistic: scaling for compute and sample-efficient continuous control

### v1

@@ Abstract
Sample efficiency in Reinforcement Learning (RL) has traditionally been driven by algorithmic enhancements. In this work, we demonstrate that scaling can also lead to substantial improvements. We conduct a thorough investigation into the interplay of scaling model capacity and domain-specific RL enhancements. These empirical findings inform the design choices underlying our proposed BRO (Bigger, Regularized, Optimistic) algorithm. The key innovation behind BRO is that strong regularization allows for effective scaling of the critic networks, which, paired with optimistic exploration, leads to superior performance. BRO achieves state-of-the-art results, significantly outperforming the leading model-based and model-free algorithms across 40 complex tasks from the DeepMind Control, MetaWorld, and MyoSuite benchmarks. BRO is the first model-free algorithm to achieve near-optimal policies in the notoriously challenging Dog and Humanoid tasks.
@@ 1 Introduction
Deep learning has seen remarkable advancements in recent years, driven primarily by the development of large neural network models (Devlin et al., 2019; Tan & Le, 2019; Dosovitskiy et al., 2020). These advancements have significantly benefited fields like natural language processing and computer vision and have been percolating to RL as well (Padalkar et al., 2023; Zitkovich et al., 2023). Interestingly, some recent work has shown that the model scaling can be repurposed to achieve sample-efficiency in discrete control (Schwarzer et al., 2023; Obando-Ceron et al., 2024), but these approaches cannot be directly translated to continuous action RL. As such, they rely on discrete action representation, whereas many physical control tasks have continuous, real valued action spaces.
@@
Conventional practice in continuous deep RL has relied on small network architectures (Haarnoja et al., 2018; Hiraoka et al., 2021; Raffin et al., 2021; D’Oro et al., 2022), with the primary focus on algorithmic improvements. These enhancements aim to achieve better sample efficiency and address key challenges such as value overestimation (Fujimoto et al., 2018; Moskovitz et al., 2021; Cetin & Celiktutan, 2023), exploration (Chen et al., 2017; Ciosek et al., 2019; Nauman & Cygan, 2023), and increasing the number of gradient steps (Nikishin et al., 2022; D’Oro et al., 2022). Additionally, evidence suggests that naive model capacity scaling can degrade performance (Andrychowicz et al., 2021; Bjorck et al., 2021). We challenge this status quo by posing a critical question: Can significant performance improvements in continuous control be achieved by combining parameter and replay scaling with existing algorithmic improvements?
In this work, we answer this question affirmatively, identifying components essential to successful scaling. Our findings are based on a thorough evaluation of a broad range of design choices, which include batch size (Obando Ceron et al., 2024), distributional Q-values techniques (Bellemare et al., 2017; Dabney et al., 2018), neural network regularizations (Bjorck et al., 2021; Nauman et al., 2024), and optimistic exploration (Moskovitz et al., 2021; Nauman & Cygan, 2023). Moreover, we carefully investigate the benefits and computational costs stemming from scaling along two axes: the number of parameters and the number of gradient steps. Importantly, we find that the former can lead to more significant performance gains while being more computationally efficient in parallelized setups.
Our work culminates in developing the BRO (Bigger, Regularized, Optimistic) algorithm, a novel sample-efficient model-free approach. BRO significantly outperforms existing model-free and model-based approaches on 40 demanding tasks from the DeepMind Control, MetaWorld, and MyoSuite benchmarks, as illustrated in Figures 1 and 2. Notably, BRO is the first model-free algorithm to achieve near-optimal performance in challenging Dog and Humanoid tasks while being 2.5 times more sample-efficient than the leading model-based algorithm, TD-MPC2. The key BRO innovation is pairing strong regularization with critic model scaling, which, coupled with optimistic exploration, leads to superior performance.
We summarize our contributions:
@@
  • BRO algorithm - we introduce the BRO algorithm, a novel model-free approach that combines regularized BroNet architecture for critic scaling with domain-specific RL enhancements. BRO achieves state-of-the-art performance on $40$ challenging tasks across diverse domains.
  • Scaling & regularization - we offer a number of insights, with the most important being: regularized critic scaling outperforms replay ratio scaling in terms of performance and computational efficiency; the inductive biases introduced by domain-specific RL improvements can be largely substituted by critic scaling, leading to simpler algorithms.
@@ 2.1 Experimental setup
We compare BRO against a variety of baseline algorithms. Firstly, we consider TD-MPC2 (Hansen et al., 2023), a model-based state-of-the-art that was shown to reliably solve the complex dog domains. Secondly, we consider SR-SAC (D’Oro et al., 2022), a sample-efficient SAC implementation that uses a large replay ratio of 32 and full-parameter resets. For completeness, we also consider CrossQ (Bhatt et al., 2023), a compute-efficient method that was shown to outperform ensemble approaches, as well as standard SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018). We run all algorithms with $10$ random seeds, except for TD-MPC2, for which we use the results provided by the original manuscript (Hansen et al., 2023). We describe the process of hyperparameter selection for all considered algorithms in Appendix E, and share BRO pseudocode in Appendix (Pseudocode 2). We implement BRO based on the JaxRL (Kostrikov, 2021) and make the code available under the following link:
@@
We consider a wide range of control tasks, encompassing a total of 40 diverse, complex continuous control tasks spanning three simulation domains: DeepMind Control (Tassa et al., 2018), MetaWorld (Yu et al., 2020), and MyoSuite (Caggiano et al., 2022) (a detailed list of environments can be found in Appendix B.2). These tasks include high-dimensional state and action spaces (with $|S|$ and $|A|$ reaching 223 and 39 dimensions), sparse rewards, complex locomotion tasks, and physiologically accurate musculoskeletal motor control. We run the algorithms for $1M$ environment steps and report the final performance unless explicitly stated otherwise. We calculate the interquartile means and confidence intervals using the RLiable package (Agarwal et al., 2021).
@@ 2.2 BRO outline and design choices
  • Optimistic – BRO uses dual policy optimistic exploration (Nauman & Cygan, 2023) and non-pessimistic (Nauman et al., 2024) quantile $Q$-value approximation (Dabney et al., 2018; Moskovitz et al., 2021) for balancing exploration and exploitation.
@@
The full details of the algorithm, along with the pseudo-code, are provided in Appendix B.1. Figure 6 summarizes the impact of removing components of BRO. We observe the biggest impact of scaling the critic capacity (scale) and replay ratio ($RR$), as well as using non-pessimistic $Q$-value, i.e. removing Clipped Double $Q$-learning (CDQ).
@@
The key contribution of this paper is showing how to enable scaling the critic network. We recall that naively increasing the critic capacity does not necessarily lead to performance improvements and that successful scaling depends on a carefully chosen suite of regularization techniques (Bjorck et al., 2021). Figure 4 shows our BroNet architecture, which, up to our knowledge, did not exist previously in the literature. The architecture begins with a dense layer followed by Layer Norm (Ba et al., 2016) and ReLU activation. Subsequently, the network comprises ResNet blocks, each consisting of two dense layers regularized with Layer Norm. Consequently, the ResNet resembles the FFN sub-layer utilized in modern LLM architectures (Xiong et al., 2020), differing primarily in the placement of the Layer Norms. Crucially, we find that BroNet scales more effectively than other architectures (Figure 3 (left)). However, the right choice of architecture and scaling is not a silver bullet. Figure 3 (right) shows that when these are plugged into the standard SAC algorithm naively, the performance is weak. The important elements are additional regularization (weight decay and network resets) and optimistic exploration (see below). Interestingly, we did not find benefits from scaling the actor networks, further discussed in Section 3.
@@
Figure 5: In order to account for sample efficiency, we report the performance averaged at $250k$, $500k$, $750k$, and $1M$ environment steps across different 5 replay ratios and 5 critic model sizes. All agents were evaluated in tasks listed in Table 3, and 10 random seeds per variant. The left figure shows performance scaling with increasing replay ratios (shapes) and model sizes (colors). The right figure examines the tradeoff between performance and computational cost when scaling replay ratios versus critic model sizes. Increasing model size leads to substantial performance improvements at lower compute costs compared to increasing the replay ratio. We present more scaling results in Appendix D, including description of model sizes in Table 5.
@@
The initial mechanism involves deactivating Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018), a commonly employed technique in reinforcement learning aimed at mitigating Q-value overestimation. For further clarification, refer to Appendix B.1.6, particularly Eq. 8 where we take the ensemble mean instead of minimum for Q-value calculation. This is surprising, perhaps, as it goes against conventional wisdom. However, some recent work has already suggested that regularization might effectively combat the overestimation (Nauman et al., 2024). We observe a much stronger effect. In Figures 6 & 8, we compare the performance of BRO with BRO that uses CDQ. This analysis indicates that using risk-neutral Q-value approximation in the presence of network regularization unlocks significant performance improvements without value overestimation (Table 1).
The second mechanism is optimistic exploration. We implement the dual actor setup (Nauman & Cygan, 2023), which employs separate policies for exploration and temporal difference updates. The exploration policy follows an optimistic upper-bound Q-value approximation, which has been shown to improve the sample efficiency of SAC-based agents (Ciosek et al., 2019; Moskovitz et al., 2021; Nauman & Cygan, 2023). In particular, we optimize the optimistic actor towards a KL-regularized Q-value upper-bound (Nauman & Cygan, 2023), with Q-value upper-bound calculated with respect to epistemic uncertainty calculated according to the methodology presented in Moskovitz et al. (2021). As shown in Figure 6, using dual actor optimistic exploration yields around $10\%$ performance improvement in BRO model.
@@ 3 Analysis
This section summarizes the results of 15,000 experiments, detailed in Table 4, which led us to develop the BRO algorithms. These experiments also provided numerous insights that we believe will be of interest to the community. We adhered to the experimental setup described in Section 2.1.
@@
Using target networks doubles the memory costs (Bhatt et al., 2023; Schwarzer et al., 2020), which can be a significant burden for large models. In Figure 8, we compare the performance of standard BRO and BRO (Fast) agents against their versions without target networks. Consistent with established understanding, we find that using target networks yields a small but significant performance improvement. However, we observe substantial variation in these effects among benchmarks and specific environments (Figure 8 & Figure 16(c)). For example, the majority of performance improvements in DMC and MS environments are attributable to specific tasks.
@@
By breaking down the results from Figure 3 into individual environments, the BroNet architecture achieves better performance in all of them, but the differences are most pronounced in the Dog environments. Therefore, we deepened our analysis with extra metrics to understand these discrepancies better. Table 1 demonstrates that BroNet outperforms the other architectures regarding final performance. The Vanilla MLP exhibits instabilities across all measured metrics, including gradient norm, overestimation, and TD error. While using the Spectral architecture maintains moderate gradient norms and overestimation, it struggles significantly with minimizing the TD error.
In Nauman et al. (2024), the authors indicate that the gradient norm and overestimation are strong indicators of poor performance in Dog environments. However, these results suggest that identifying a single cause for the challenges in training a reinforcement learning agent is difficult, highlighting the complexity of these systems and the multifaceted nature of their performance issues.
@@
As illustrated in Figure 8(b), even complex tasks like Dog Walk or Dog Trot can be reliably solved by combining existing algorithmic improvements with critic model scaling within 1 million environment steps. However, some tasks remain unsolved within this limit (e.g., Humanoid Run or Acrobot Swingup). Tailoring algorithms to single tasks risks overfitting to specific issues. Therefore, we advocate for standardized benchmarks that reflect the sample efficiency of modern algorithms. This standardization would facilitate consistent comparison of approaches, accelerate advancements by focusing on a common set of challenging tasks, and promote the development of more robust and generalizable RL algorithms. On that note, in Appendix G, we report BRO performance at earlier stages of the training.
@@ 4 Related Work
Sample efficiency through algorithmic improvements
A significant effort in RL has focused on algorithmic improvements. One recurring theme is controlling value overestimation (Fujimoto et al., 2018; Moskovitz et al., 2021; Cetin & Celiktutan, 2023). For instance, Fujimoto et al. (2018) proposed Clipped Double Q-learning (CDQ), which updates policy and value networks using a lower-bound Q-value approximation. However, since a pessimistic lower-bound can slow down the learning, Moskovitz et al. (2021) introduced approach that tunes pessimism online. Recently, Nauman et al. (2024) showed that layer normalization can improve performance without value overestimation, eliminating the need for pessimistic Q-learning. A notable effort has also focused on optimistic exploration (Wang et al., 2020; Moskovitz et al., 2021). Various methods have been developed to increase sample efficiency via exploration that is greedy with respect to a Q-value upper bound. These include closed-form transformations of the pessimistic policy (Ciosek et al., 2019) or using a dual actor network dedicated to exploration (Nauman & Cygan, 2023).
Sample efficiency through scaling
Recent studies demonstrated the benefits of model scaling when pre-training on large datasets (Driess et al., 2023; Schubert et al., 2023; Taiga et al., 2023) or in pure offline RL setups (Kumar et al., 2023). Additionally, model scaling has proven advantageous for model-based online RL (Hafner et al., 2023; Hansen et al., 2023; Wang et al., 2024). However, in these approaches, most of the model scale is dedicated to world models, leaving the value network relatively small. Notably, Schwarzer et al. (2023) found that increasing the scale of the encoder network improves performance for DQN agents, but did not study increasing the capacity of the value network. Various studies indicate that naive scaling of the value model leads to performance deterioration (Bjorck et al., 2021; Obando-Ceron et al., 2024; Farebrother et al., 2024). For example, Bjorck et al. (2021) demonstrated that spectral normalization enables stable training with relatively large ResNet architectures, yielding performance improvements. In addition to model size scaling, the community has investigated the effectiveness of replay ratio scaling (i.e., increasing the number of gradient steps for every environment step) (Hiraoka et al., 2021; Nikishin et al., 2022; Li et al., 2022). Recent works have shown that a high replay ratio can improve performance across various algorithms, provided the neural networks are regularized (Li et al., 2022; D’Oro et al., 2022). In this context, layer normalization and full-parameter resets have been particularly effective (Schwarzer et al., 2023; Lyle et al., 2024; Nauman et al., 2024).

### v2

@@ Abstract
Sample efficiency in Reinforcement Learning (RL) has traditionally been driven by algorithmic enhancements. In this work, we demonstrate that scaling can also lead to substantial improvements. We conduct a thorough investigation into the interplay of scaling model capacity and domain-specific RL enhancements. These empirical findings inform the design choices underlying our proposed BRO (Bigger, Regularized, Optimistic) algorithm. The key insight behind BRO is that strong regularization allows for effective scaling of the critic networks, which, paired with optimistic exploration, leads to superior performance. BRO achieves state-of-the-art results, significantly outperforming the leading model-based and model-free algorithms across 40 complex tasks from the DeepMind Control, MetaWorld, and MyoSuite benchmarks. BRO is the first model-free algorithm to achieve near-optimal policies in the notoriously challenging Dog and Humanoid tasks.
@@ 1 Introduction
11footnotetext: Ideas NCBR; 2University of Warsaw, 3Warsaw University of Technology, 4Polish Academy of Sciences, 5Nomagic. Correspondence to: Michal Nauman <nauman.mic@gmail.com>
Deep learning has seen remarkable advancements in recent years, driven primarily by the development of large neural network models (Devlin et al., 2019; Tan & Le, 2019; Dosovitskiy et al., 2020). These advancements have significantly benefited fields like natural language processing and computer vision and have been percolating to RL as well (Padalkar et al., 2023; Zitkovich et al., 2023). Interestingly, some recent work has shown that the model scaling can be repurposed to achieve sample efficiency in discrete control (Schwarzer et al., 2023; Obando-Ceron et al., 2024), but these approaches cannot be directly translated to continuous action RL. As such, they rely on discrete action representation, whereas many physical control tasks have continuous, real-valued action spaces.
@@
Conventional practice in continuous deep RL has relied on small network architectures (Haarnoja et al., 2018; Hiraoka et al., 2021; Raffin et al., 2021; D’Oro et al., 2022), with the primary focus on algorithmic improvements. These enhancements aim to achieve better sample efficiency and address key challenges such as value overestimation (Fujimoto et al., 2018; Moskovitz et al., 2021; Cetin & Celiktutan, 2023), exploration (Chen et al., 2017; Ciosek et al., 2019; Nauman & Cygan, 2023a), and increasing the number of gradient steps (Nikishin et al., 2022; D’Oro et al., 2022). Additionally, evidence suggests that naive model capacity scaling can degrade performance (Andrychowicz et al., 2021; Bjorck et al., 2021). We challenge this status quo by posing a critical question: Can significant performance improvements in continuous control be achieved by combining parameter and replay scaling with existing algorithmic improvements?
In this work, we answer this question affirmatively, identifying components essential to successful scaling. Our findings are based on a thorough evaluation of a broad range of design choices, which include batch size (Obando Ceron et al., 2024), distributional Q-values techniques (Bellemare et al., 2017; Dabney et al., 2018), neural network regularizations (Bjorck et al., 2021; Nauman et al., 2024), and optimistic exploration (Moskovitz et al., 2021; Nauman & Cygan, 2023a). Moreover, we carefully investigate the benefits and computational costs stemming from scaling along two axes: the number of parameters and the number of gradient steps. Importantly, we find that the former can lead to more significant performance gains while being more computationally efficient in parallelized setups.
Our work culminates in developing the BRO (Bigger, Regularized, Optimistic) algorithm, a novel sample-efficient model-free approach. BRO significantly outperforms existing model-free and model-based approaches on 40 demanding tasks from the DeepMind Control, MetaWorld, and MyoSuite benchmarks, as illustrated in Figures 1 and 2. Notably, BRO is the first model-free algorithm to achieve near-optimal performance in challenging Dog and Humanoid tasks while being 2.5 times more sample-efficient than the leading model-based algorithm, TD-MPC2. The key BRO innovation is pairing strong regularization with critic model scaling, which, coupled with optimistic exploration, leads to superior performance. We summarize our contributions:
@@
  • BroNet architecture & BRO algorithm - we introduce the BRO algorithm, a novel model-free approach that combines BroNet architecture for critic scaling with domain-specific RL enhancements. BRO achieves state-of-the-art performance on $40$ challenging tasks across diverse domains.
  • Scaling & regularization - we offer several insights, with the most important being: regularized critic scaling outperforms replay ratio scaling in terms of performance and computational efficiency; the inductive biases introduced by domain-specific RL improvements can be largely substituted by critic scaling, leading to simpler algorithms.
@@ 2.1 Experimental setup
We compare BRO against a variety of baseline algorithms. Firstly, we consider TD-MPC2 (Hansen et al., 2023), a model-based state-of-the-art that was shown to reliably solve the complex dog domains. Secondly, we consider SR-SAC (D’Oro et al., 2022), a sample-efficient SAC implementation that uses a large replay ratio of 32 and full-parameter resets. For completeness, we also consider CrossQ (Bhatt et al., 2023), a compute-efficient method that was shown to outperform ensemble approaches, as well as standard SAC (Haarnoja et al., 2018) and TD3 (Fujimoto et al., 2018). We run all algorithms with $10$ random seeds, except for TD-MPC2, for which we use the results provided by the original manuscript (Hansen et al., 2023). We describe the process of hyperparameter selection for all considered algorithms in Appendix D, and share BRO pseudocode in Appendix (Pseudocode 2). We implement BRO based on the JaxRL (Kostrikov, 2021) and make the code available under the following link: https://github.com/naumix/BiggerRegularizedOptimistic
@@
We consider a wide range of control tasks, encompassing a total of 40 diverse, complex continuous control tasks spanning three simulation domains: DeepMind Control (Tassa et al., 2018), MetaWorld (Yu et al., 2020), and MyoSuite (Caggiano et al., 2022) (a detailed list of environments can be found in Appendix C). These tasks include high-dimensional state and action spaces (with $|S|$ and $|A|$ reaching 223 and 39 dimensions), sparse rewards, complex locomotion tasks, and physiologically accurate musculoskeletal motor control. We run the algorithms for $1M$ environment steps and report the final performance unless explicitly stated otherwise. We calculate the interquartile means and confidence intervals using the RLiable package (Agarwal et al., 2021).
(a) DeepMind Control (DMC)
(b) MetaWorld (MW)
(c) MyoSuite (MS)
@@ 2.2 BRO outline and design choices
  • Optimistic – BRO uses dual policy optimistic exploration (Nauman & Cygan, 2023a) and non-pessimistic (Nauman et al., 2024) quantile $Q$-value approximation (Dabney et al., 2018; Moskovitz et al., 2021) for balancing exploration and exploitation.
@@
The full details of the algorithm, along with the pseudo-code, are provided in Appendix B. Figure 7 summarizes the impact of removing components of BRO. We observe the biggest impact of scaling the critic capacity (scale) and replay ratio ($RR$), as well as using non-pessimistic $Q$-value, i.e. removing Clipped Double $Q$-learning (CDQ).
@@
The key contribution of this paper is showing how to enable scaling the critic network. We recall that naively increasing the critic capacity does not necessarily lead to performance improvements and that successful scaling depends on a carefully chosen suite of regularization techniques (Bjorck et al., 2021). Figure 5 shows our BroNet architecture, which, up to our knowledge, did not exist previously in the literature. The architecture begins with a dense layer followed by Layer Norm (Ba et al., 2016) and ReLU activation. Subsequently, the network comprises ResNet blocks, each consisting of two dense layers regularized with Layer Norm. Consequently, the ResNet resembles the FFN sub-layer utilized in modern LLM architectures (Xiong et al., 2020), differing primarily in the placement of the Layer Norms. Crucially, we find that BroNet scales more effectively than other architectures (Figure 3(c) (left)). However, the right choice of architecture and scaling is not a silver bullet. Figure 3(c) (right) shows that when these are plugged into the standard SAC algorithm naively, the performance is weak. The important elements are additional regularization (weight decay and network resets) and optimistic exploration (see below). Interestingly, we did not find benefits from scaling the actor networks, further discussed in Section 3.
@@
Figure 6: To account for sample efficiency, we report the performance averaged at $250k$, $500k$, $750k$, and $1M$ environment steps across different 5 replay ratios and 5 critic model sizes. All agents were evaluated in tasks listed in Table 3, and 10 random seeds per variant. The left figure shows performance scaling with increasing replay ratios (shapes) and model sizes (colors). The right figure examines the tradeoff between performance and computational cost when scaling replay ratios versus critic model sizes. Increasing model size leads to substantial performance improvements at lower compute costs compared to increasing the replay ratio. We present more scaling results in Appendix E, including a description of model sizes in Table 7.
@@
The initial mechanism involves deactivating Clipped Double Q-learning (CDQ) (Fujimoto et al., 2018), a commonly employed technique in reinforcement learning aimed at mitigating Q-value overestimation. For further clarification, refer to Appendix B.6, particularly Eq. 8 where we take the ensemble mean instead of minimum for Q-value calculation. This is surprising, perhaps, as it goes against conventional wisdom. However, some recent work has already suggested that regularization might effectively combat the overestimation (Nauman et al., 2024). We observe a much stronger effect. In Figures 7 & 9, we compare the performance of BRO with BRO that uses CDQ. This analysis indicates that using risk-neutral Q-value approximation in the presence of network regularization unlocks significant performance improvements without value overestimation (Table 1).
The second mechanism is optimistic exploration. We implement the dual actor setup (Nauman & Cygan, 2023a), which employs separate policies for exploration and temporal difference updates. The exploration policy follows an optimistic upper-bound Q-value approximation, which has been shown to improve the sample efficiency of SAC-based agents (Ciosek et al., 2019; Moskovitz et al., 2021; Nauman & Cygan, 2023a). In particular, we optimize the optimistic actor towards a KL-regularized Q-value upper-bound (Nauman & Cygan, 2023a), with Q-value upper-bound calculated with respect to epistemic uncertainty calculated according to the methodology presented in Moskovitz et al. (2021). As shown in Figure 7, using dual actor optimistic exploration yields around $10\%$ performance improvement in the BRO model.
@@ 3 Analysis
This section summarizes the results of 15,000 experiments, detailed in Table 2, which led us to develop the BRO algorithms. These experiments also provided numerous insights that we believe will be of interest to the community. We adhered to the experimental setup described in Section 2.1. We also present additional experimental results in Appendix E.
@@
Using target networks doubles the memory costs (Schwarzer et al., 2020; Bhatt et al., 2023; Lee et al., 2024), which can be a significant burden for large models. In Figure 9, we compare the performance of standard BRO and BRO (Fast) agents against their versions without target networks. Consistent with established understanding, we find that using target networks yields a small but significant performance improvement. However, we observe substantial variation in these effects among benchmarks and specific environments (Figure 9 & Figure 16). For example, the majority of performance improvements in DMC and MS environments are attributable to specific tasks.
@@
By breaking down the results from Figure 3(c) into individual environments, the BroNet architecture achieves better performance in all of them, but the differences are most pronounced in the Dog environments. Therefore, we deepened our analysis with extra metrics to understand these discrepancies better. Table 1 demonstrates that BroNet outperforms the other architectures regarding final performance. The Vanilla MLP exhibits instabilities across all measured metrics, including gradient norm, overestimation, and TD error. While using the Spectral architecture maintains moderate gradient norms and overestimation, it struggles significantly with minimizing the TD error.
In (Nauman et al., 2024), the authors indicate that the gradient norm and overestimation are strong indicators of poor performance in Dog environments. However, these results suggest that identifying a single cause for the challenges in training a reinforcement learning agent is difficult, highlighting the complexity of these systems and the multifaceted nature of their performance issues.
@@
As illustrated in Figure 9(b), even complex tasks like Dog Walk or Dog Trot can be reliably solved by combining existing algorithmic improvements with critic model scaling within 1 million environment steps. However, some tasks remain unsolved within this limit (e.g., Humanoid Run or Acrobot Swingup). Tailoring algorithms to single tasks risks overfitting to specific issues. Therefore, we advocate for standardized benchmarks that reflect the sample efficiency of modern algorithms. This standardization would facilitate consistent comparison of approaches, accelerate advancements by focusing on a common set of challenging tasks, and promote the development of more robust and generalizable RL algorithms. On that note, in Appendix F, we report BRO performance at earlier stages of the training.
BroNet architecture is useful beyond continuous control
We design additional experiments to test the effectiveness of the naive application of BroNet to popular offline reinforcement learning problems in two offline RL benchmarks: AntMaze (6 tasks); and Adroit (9 tasks) (Fu et al., 2020). We run Behavioral Cloning (BC) in pure offline (Sutton & Barto, 2018), Implicit Q-Learning (IQL) offline + fine-tuning (Kostrikov et al., 2022), as well as online reinforcement learning with offline data (Ball et al., 2023). We run all these algorithms with the default MLP network, as well as BroNet backbone. As shown in Figure 11, we find that the naive application of BroNet leads to performance improvements across all tested algorithms.
Figure 11: We test three scenarios: offline (comparing vanilla BC to BroNet-based BC), offline fine-tuning (comparing vanilla IQL to BroNet-based IQL), and online with offline data (comparing vanilla SAC to BroNet-based SAC). The solid line represents BRO-based and the dashed line represents vanilla variants. Negative values on the X-axis refer to offline training. 10 seeds per task.
@@ 4 Related Work
4.1 Sample efficiency through algorithmic improvements
A significant effort in RL has focused on algorithmic improvements. One recurring theme is controlling value overestimation (Fujimoto et al., 2018; Moskovitz et al., 2021; Cetin & Celiktutan, 2023). For instance, Fujimoto et al. (2018) proposed Clipped Double Q-learning (CDQ), which updates policy and value networks using a lower-bound Q-value approximation. However, since a pessimistic lower-bound can slow down learning, Moskovitz et al. (2021) introduced an approach that tunes pessimism online. Recently, Nauman et al. (2024) showed that layer normalization can improve performance without value overestimation, eliminating the need for pessimistic Q-learning.
A notable effort has also focused on optimistic exploration (Wang et al., 2020; Moskovitz et al., 2021). Various methods have been developed to increase sample efficiency via exploration that is greedy with respect to a Q-value upper bound. These include closed-form transformations of the pessimistic policy (Ciosek et al., 2019) or using a dual actor network dedicated to exploration (Nauman & Cygan, 2023a).
4.2 Sample efficiency through scaling
Recent studies demonstrated the benefits of model scaling when pre-training on large datasets (Driess et al., 2023; Schubert et al., 2023; Taiga et al., 2023) or in pure offline RL setups (Kumar et al., 2023). Additionally, model scaling has proven advantageous for model-based online RL (Hafner et al., 2023; Hansen et al., 2023; Wang et al., 2024). However, in these approaches, most of the model scale is dedicated to world models, leaving the value network small. Notably, Schwarzer et al. (2023) found that increasing the scale of the encoder network improves performance for DQN agents, but did not study increasing the capacity of the value network. Various studies indicate that naive scaling of the value model leads to performance deterioration (Bjorck et al., 2021; Obando-Ceron et al., 2024; Farebrother et al., 2024). For example, Bjorck et al. (2021) demonstrated that spectral normalization enables stable training with relatively large ResNets with performance improvements.
In addition to model size scaling, the community has investigated the effectiveness of replay ratio scaling (i.e., increasing the number of gradient steps for every environment step) (Hiraoka et al., 2021; Nikishin et al., 2022; Li et al., 2022). Recent works have shown that a high replay ratio can improve performance across various algorithms in both continuous and discrete MDPs, provided the neural networks are regularized (Li et al., 2022; D’Oro et al., 2022). In this context, layer normalization and full-parameter resets have been particularly effective (Schwarzer et al., 2023; Lyle et al., 2024; Nauman et al., 2024).